# Error Correction in Dynamical Codes

### v1

@@ 1 Introduction
Floquet codes generalize "static" stabilizer codes, which in this paper will refer to codes that are defined by either a fixed gauge group or a fixed stabilizer group. Introduced by Hastings and Haah as the first example of Floquet code, the honeycomb code [1][2] consists of rounds of checks that anti-commutes with the previous rounds, creating a periodic change in the stabilizer generators and the logical operators. The stabilizer group at each time step is called the instantaneous stabilizer group (ISG). To decode, the checks from multiple rounds are combined to obtain syndrome information. One advantage of the measurement scheme is that only low weight measurements are needed, and they combine to give syndrome information of high weight stabilizers. Measurement errors also become important to recover the logical information, because the logical operators are updated by some measurements to commute with the next ISG.
@@
Since the first paper on Floquet codes, numerous additional examples of Floquet codes have been found. This includes planar Floquet codes [3][2][4][5], 2D Floquet codes defined on hyperbolic surfaces[6][7], and 3D Floquet codes[8][9][10]. There are also CSS version of Floquet code and codes with non-periodicity [10][4]. The threshold and performance of Floquet code have been demonstrated in [11][12]. Progress has also been made in the theoretical understanding of Floquet codes, including results based on measurement quantum cellular automata [13], adiabatic paths of gapped Hamiltonians [14], and unitary loops [15].
@@
Generalizing and broadening the idea of Floquet codes, we wish to understand the framework of quantum codes that are defined by a sequence of measurements. Our study is in part driven by insights on code deformation and code switching [16][17] in implementing a universal set of gates: For example, code switching between Reed Muller code and Steane code enables the transversal implementation of universal logical gates [18]. As a dynamical code consists of several rounds of ISG, it can potentially circumvent the no-go theorem by Eastin and Knill [19] which shows that transversal operators on any non-trivial quantum code belong to a finite group, and thus cannot be universal.
We use the term dynamical code instead of Floquet code, as we do not exclude the possibility of a non-periodic sequence of Pauli measurements. Most previous works of Floquet codes have concentrated on codes defined on three-colorable graphs, here we look at more general codes. In this paper, a dynamical code is more generally defined as a code that involves rounds of Pauli measurements. Each round of measurements consists of only commuting Pauli operators. In particular, a two round dynamical code is a subsystem code, such as the Bacon Shor code [20], which consists of geometrically local gauges and non-local stabilizers.
@@
The concept of distance is well understood for "static" stabilizer codes but more ambiguous for dynamical codes. The instantaneous stabilizer groups of a dynamical code inherit the same concept of distance as that of a static code. However, the same cannot be said for the distance of a dynamical code, as a measurement scheme can irreversibly remove some syndrome information from the code while other parts of the syndromes are revealed. Previous work by Gottesman [21] introduces the notion of masking. A permanently masked element is defined as a stabilizer whose eigenvalue cannot be obtained again in the future. These are elements that have been irreversibly removed without revealing their syndrome information. Masking also has practical relevance for partial syndrome measurements[22].
Here we present a general method to determine the set of all the syndrome information that can be learnt from a sequence of measurements and determine the unmasked distance of a dynamical code, with the assumption that the set of syndromes for error correction is obtained only through subsequent rounds of measurements. We named this the distance algorithm.
This work is related to the algorithms proposed by Delfosse and Paetznick[23]. There the authors proposed the outcome code for a circuit with both measurements and unitary gates. They also constructed a spacetime code to correct for circuit faults. The syndrome information that we obtained here coincides with the outcome code. Our work goes beyond this and further classifies the syndrome information and defines the distance for dynamical codes. We determine which round(s) of the ISG are the syndrome information for, and we can determine the time window one requires to obtain enough syndrome information for a given round. Our algorithm further gives the set of permanently masked stabilizers, whose syndromes one can never obtain, and the temporarily masked stabilizers, whose syndromes can be obtained by inserting new measurements into existing measurement sequence. Based on this, we formulate the unmasked distance for each ISG and find an upper bound for the distance of a dynamical code.
@@
We have considered only dynamical codes, which can be seen as a special case of Clifford circuits where the unitaries between measurements are set to identity. To be able to correct for circuit faults, it is a necessary but not sufficient condition that when the unitaries are set to identity, the code (now a dynamical code) has good error correcting properties. Our algorithm can easily be generalized to include unitary gates by pushing all the unitaries to the end of the circuit.
Further, we simplified the problem by considering the scenario where the errors only occur in round 0 of the dynamical code. This allows us to obtain perfect syndrome information from later rounds and give an upper bound to the error correcting capacity of the dynamical code.
It is worth noting that this bound assumes perfect measurements. However, in the case of measurement errors, it is possible to create a smaller set of errors in space-time that will lead to a logical error[3]. Our results lack generalization to the fault tolerant context, as decoding in this context proves significantly more challenging than the static code case. This stems from the fact that errors occurring at later rounds and measurement errors have the potential to change the syndrome bits from earlier rounds. Consequently, understanding the desired properties for a dynamical code to have a threshold is notably more intricate. The challenges posed by decoding in the presence of measurement errors and errors in future rounds underscore the complexities associated with extending the analysis to the fault-tolerant regime. However, our results can contribute to a better understanding of dynamical codes by establishing the foundation for the code properties that are inherent to dynamical codes. We hope that the results here can later be extended to analyze the fault-tolerance aspect of dynamical codes.
@@
For a quantum code with a fixed set of stabilizer generators, the problem of finding the distance and fault-tolerant gates has been previously studied from the angle of embedding them in Euclidean space, as locality is very important in practical code design. Bravyi and Terhal [24] proved an upper bound on local code distance: $d=O(L^{D-1})$. Further, Bravyi and König [25] showed that a $D$-dimensional local code can only transversally implement encoded gates from the set $\mathcal{C}^{(D)}$, with $\mathcal{C}^{(1)}$ the Pauli group and $\mathcal{C}^{(j)}$ as the $j^{\mathrm{th}}$ level of the Clifford hierarchy. Thus, we are interested in answering this question: To what extent does these locality restrictions apply to geometrically local dynamical codes?
@@
We generalize Pastawski and Yoshida’s results [26] and give a no-go theorem for 2D geometrically local dynamical code to support transversal non-Clifford gates even with some amount of long-range connectivity and non-local stabilizers. This result also extends to higher dimensions. Specifically, we prove that if a logical qubit $q_{L}$ can be supported on a region sufficiently far away from the qubits with long range connectivity, then a transversal gate that implements a logical single qubit gate on $q_{L}$ must be from the $D^{\mathrm{th}}$ level Clifford hierarchy. Furthermore, if there is limited amount of long range connectivity, then the code cannot support a logical gate from a higher Clifford hierarchy.
The paper is structured as follows. In Section 2, we give a brief introduction to quantum error correcting codes and define some notations for the rest of the paper. In Section 3, we define distance for dynamical codes and discuss the logical operators and syndrome information in the presence of measurements. In Subsection 4.1, we give an intuition through examples to demonstrate that it can be tricky to find the distance of a given round in the dynamical code. In Subsection 4.3, we present the algorithm for finding the unmasked distance of dynamical codes, and show that the algorithm can output all the unmasked stabilizers and permanently masked stabilizers of an ISG. In Section 5, we apply the tools developed for the distance algorithm to analyze the initialization and masking properties of Floquet codes. In Section 6, we generalize previous works for geometrically local stabilizer codes to dynamical codes and also present a no-go theorem for non-Clifford transversal gates on 2D geometrically local dynamical codes.
@@ 2.1 Stabilizer code
A stabilizer code $\mathcal{Q}=\mathcal{Q}(S)$ with parameters $[[n,k,d]]$ is a code defined on the Hilbert space $\mathcal{H}$ of $n$ qubits. It has $n-k$ Pauli stabilizer generators $S$ and $k$ logical qubits. The codeword space $\mathcal{C}$ is a subspace of $\mathcal{H}$, and contains the $+1$ eigenvectors for stabilizers in $S$. The non-trivial logical representatives are given by the set $N(S)\backslash\langle S\rangle$, where $N(S)$ is the normalizer of $\langle S\rangle$ defined as $N(S)=\{p\in\mathcal{P}_{n}:\langle S\rangle p=p\langle S\rangle\},\mathcal{P}_{n}$ is the $n$ qubit Pauli group, with $\mathcal{P}=\{\pm 1,\pm i,I,X,Y,Z\}$, $\mathcal{P}_{n}=\mathcal{P}^{\otimes n}$ and $\langle S\rangle$ is the group generated by $S$. $d$ is the distance, which for a stabilizer code is the smallest weight Pauli operator $L$ for which $L$ commutes with all elements of $S$ but is not itself in $\langle S\rangle$. The distance is the lowest weight of elements in $N(S)\backslash\langle S\rangle$.
@@ 2.4 Update rules
Let $\mathcal{S}$ be the stabilizer generators with a stabilizer state |$\ket{\psi}$ being either in $+1$ or $-1$ eigenstate of the generators.
Let $m$ be a Pauli measurement performed on |$\ket{\psi}$, and denote the outcome of $m$ by $O(m)\in\{\pm 1\}$.
@@
    If ±$\pm m\in\langle\mathcal{S}\rangle$, then the outcome is fixed by the eigenvalues of stabilizers for |$\ket{\psi}$, and the state remains unchanged.
@@
    If ±$\pm m\notin\langle\mathcal{S}\rangle$ and $[m,s]=0\,\forall s\in\mathcal{S}$, then we update the set of stabilizer generators: $\mathcal{S}\to\mathcal{S}\cup\{O(m)\cdot m\}$. This assumes that $m$ is not a logical operator.
@@
Let $L$ be a logical operator of a stabilizer group $\langle S\rangle$ and let |$\ket{\psi}$ be an eigenstate of $L$.
Let $m$ be a Pauli measurement performed on |$\ket{\psi}$ and denote the outcome by $O(m)\in\{\pm 1\}$.
@@
    If $m=(-1)^{a}\cdot L$, then $O(m)\cdot(-1)^{a}$ gives the eigenvalue of $L$ for the state |$\ket{\psi}$, and the logical operator remains unchanged.
@@
    If $m$ anti-commutes with $L$ and commutes with $\langle S\rangle$, then $L$ is updated to $O(m)\cdot m$. The new state is a $+1$ eigenstate of $O(m)\cdot m$ instead of $L$.
@@ 2.5 Dynamical code and Subsystem code
Given a set of independent stabilizer generators $S$, a destabilizer $\kappa\in K$ of a stabilizer generator $s\in S$ is a Pauli operator in $\mathcal{P}_{n}/\mathbb{C}$ that only anti-commutes with $s$, but commutes with $S\backslash s$ and the bare logical operators of the gauge group $\mathcal{G}=\langle S\cup K\rangle$.

### v2

@@ 1 Introduction
Floquet codes extend the framework of conventional “static” stabilizer codes–those defined by fixed stabilizer or gauge groups–by introducing time-dependent measurements. The honeycomb code introduced in [1, 2] is a first example of a new paradigm of quantum codes that features periodic sequence of measurements that anti-commute with the previous round of measurements, leading to a periodic change in both the stabilizer group and the logical representatives. At each moment, the stabilizers form what’s known as the instantaneous stabilizer group (ISG). Decoding involves combining measurement outcomes from multiple rounds. One key advantage of Floquet code is the low weight measurements, which are combined to reveal the syndrome of high-weight stabilizers—unlocking the power of error correction with minimal overhead.
@@
Since the first paper on Floquet codes, numerous additional examples of Floquet codes have been found. This includes planar Floquet codes [3, 2, 4, 5], 2D Floquet codes defined on hyperbolic surfaces[6, 7], and 3D Floquet codes[8, 9, 10]. There are also CSS version of Floquet codes and codes with non-periodicity [10, 4]. The threshold and performance of Floquet codes have been demonstrated in [11, 12]. Progress has also been made in the theoretical understanding of Floquet codes, including results based on measurement quantum cellular automata [13], adiabatic paths of gapped Hamiltonians [14], and unitary loops [15]. Later, new Floquet codes that are not based on colorable graphs have also been found, including the Floquet Bacon-Shor code [16] and Floquet Haah code [17].
@@
Generalizing and broadening the idea of Floquet codes, we wish to understand the framework of quantum codes that are defined by a sequence of measurements. Our study is in part driven by insights on code deformation and code switching [18, 19] in implementing a universal set of gates: For example, code switching between Reed Muller code and Steane code enables the transversal implementation of universal logical gates [20]. As a dynamical code consists of several rounds of ISGs, it can potentially circumvent the no-go theorem by Eastin and Knill [21] which shows that transversal operators on any non-trivial quantum code belong to a finite group, and thus cannot be universal. We use the term dynamical code instead of Floquet code, as we do not exclude the possibility of a non-periodic sequence of Pauli measurements. Most previous works of Floquet codes have concentrated on codes defined on three-colorable graphs, here we look at more general codes: A dynamical code is more generally defined as a code that involves rounds of Pauli measurements. Each round of measurements consists of only commuting Pauli operators. In particular, a two round dynamical code is a subsystem code, such as the Bacon-Shor code [22], which consists of geometrically local gauges and non-local stabilizers.
In this work, we present a comprehensive study of the framework of dynamical code using QEC language. We formalize the concept of masking with respect to a dynamical code and introduce a notion of distance that generalizes known notions of distance of stabilizer and subsystem codes. Building on this theoretical foundation, we will also introduce an efficient algorithm to determine the syndromes and therefore the distance for a given dynamical code. Making use of the structure of this algorithm, we derive new insights into the initialization of arbitrary Floquet codes. Finally, we establish fundamental limitations by extending the Bravyi–König no-go theorem to the setting of dynamical codes. See Section 2 for an overview of the main results.
The paper is structured as follows. We start with an overview of our results in Section 2. In Section 3, we give a brief introduction to quantum error correcting codes and define some notations for the rest of the paper. In Section 4, we define distance for dynamical codes and discuss the logical operators and syndrome information in the presence of measurements. Section 5 discusses the distance algorithm: In Subsection 5.1, we give an intuition through examples to demonstrate that it can be tricky to find the distance of a given round in the dynamical code. In Subsection 5.3, we present the algorithm for finding the unmasked distance of dynamical codes, and show that the algorithm can output all the unmasked stabilizers and permanently masked stabilizers of an ISG. In Section 6, we apply the tools developed for the distance algorithm to analyze the initialization and masking properties of Floquet codes. In Section 7, we generalize previous works for geometrically local stabilizer codes to dynamical codes and also present a no-go theorem for non-Clifford transversal gates on 2D geometrically local dynamical codes.
2 Overview
We now give an overview of our contributions in this work.
Framework of dynamical code
In Section 4, we extend the framework of quantum error correcting codes from stabilizer and subsystem codes to dynamical codes.
Subsection 4.1 present a tighter notion of distance for dynamical code. Our definition captures the fact that, unlike static codes, the correctability of errors in a dynamical code depends not only on the current stabilizer group but also on the syndrome information recoverable from future measurements. The distance of a dynamical code is defined as the minimum unmasked distance over all instantaneous stabilizer groups (ISGs), where the unmasked distance quantifies the set of correctable errors under the given measurement sequence.
@@ 2 Overview
What makes the previous notions of distance inadequate for dynamical codes is that a measurement scheme can irreversibly remove some syndrome information from the code while other parts of the syndromes are revealed. This is the concept of masking which has been previously introduced by Gottesman [23]. For instance, a stabilizer generator can be permanently masked if its eigenvalue cannot be obtained again in the future. These are elements that have been irreversibly removed without revealing their syndrome information.
We classify each stabilizer generator in every instantaneous stabilizer group into three categories:
  • Permanently masked stabilizers: Stabilizers whose syndromes cannot be obtained using given sequence of measurements.
  • Temporarily masked stabilizers: Stabilizers whose syndromes have yet to be obtained with the current sequence but can still be obtained if we choose further measurements carefully.
  • Unmasked stabilizers: Stabilizers whose syndromes can be obtained from outcomes of the current sequence of measurements.
Using this classification, we define the unmasked distance of an ISG in Definition 22. We prove that this notion yields a tighter upper bound on the set of correctable errors than prior definitions, by generalizing the Knill–Laflamme condition to the setting where only partial syndrome information is available. We conjecture that the unmasked distance is tight in the absence of measurement errors.
In Subsection 4.3, we provide a general theoretical framework for analyzing how logical operators and logical errors evolve over time in a dynamical code. Unlike static codes, where logical representatives are fixed, in a dynamical code the logical representatives and their values are inherently spacetime-dependent: the outcome of a logical operator can depend both on the spacetime location of errors and on the measurement outcomes of relevant stabilizers at different times.
Theorem 29 in Subsection 4.4 shows that, in addition to where the errors happen, their temporal positions crucially affect the observed syndromes. This result highlights the need for decoding algorithms that take into account not just which qubits are erroneous, but also when those errors occur, relative to the measurement sequence.
Algorithm for finding distance of a dynamical code
In Section 5, we present the distance algorithm.
The distance algorithm is an efficient classical algorithm that determines the set of all the syndrome information that can be learnt from a fixed measurement window (see Section 5). The objective is to classify the stabilizer generators of a fixed ISG into three types: unmasked, temporarily masked, and permanently masked. This classification captures which syndromes can be reliably extracted in a finite time window and provides an upper bound—the unmasked distance—on the distance of the code (Definition 22).
The algorithm proceeds by tracking the evolution of stabilizers and measurement outcomes through a set of commutation-based update rules (Subsection 5.3). Two sets, $C$ and $V$, are iteratively updated to record the transformed stabilizer generators and accumulated measurements, respectively. The unmasked stabilizers are obtained by computing the intersection $\langle C\rangle\cap\langle V\rangle$, implemented via the Zassenhaus algorithm (Algorithm 1). From this, the syndromes for unmasked stabilizers are reconstructed (Algorithm 2). The algorithm also determines permanently masked stabilizers and their destabilizers (Algorithm 3), as well as the temporarily masked ones (Algorithm 4).
Our framework provides a concrete and efficient tool to determine the partial decoding power of a dynamical code under realistic measurement constraints. The algorithm runs in $O(n^{3})$ time, with the main cost coming from the subroutine to compute generator intersections. Several illustrative examples are given in Subsection 5.4.
@@
Going beyond code construction, the distance algorithm can be extended with relative ease to obtain the outcome code in [24] since the unmasked stabilizers are essentially detector cells restricted to the time window being analyzed, and that Clifford operators can be included between measurement rounds. But beyond this, we can use the distance algorithm to obtain gauge elements of a subsystem spacetime code, since we also take into account other elements including the permanently and temporarily masked stabilizers with their destabilizers. We also track which round(s) of the ISGs the syndrome information is for, and the time window one requires to obtain enough syndrome information for a given round. Although we consider errors in round 0 and assume perfect measurements (in the presence of measurement errors, it is possible for a smaller set of errors in space-time to result in a logical error [3]), our results can contribute to a better understanding of dynamical codes by establishing some code properties that are inherent to dynamical codes. The results here can also be extended to analyze the fault-tolerance aspect of dynamical codes and then to that of spacetime codes. A better understanding of the distance for dynamical codes will be insightful for understanding the distance and fault tolerance for a Clifford circuit, even eventually extending to include more general noise models.
One aspect of our algorithm is that it identifies which stabilizer generators are temporarily masked and which are permanently masked. This is critical information for a dynamical coding protocol that is described in a modular way: Temporarily masked generators from one module may usefully contribute syndrome information into a later module whereas syndrome information from permanently masked generators will never be useful, regardless of the choice of later modules.
Lastly, we remark that there have been several related works. The distance algorithm helps complete the picture by adding anti-commuting measurements into the framework of [25, 23], which translate circuits into quantum codes. The syndrome information from the distance algorithm contains the same information as the algorithm in [24]. The connection between understanding the detector cell/spacetime code picture to designing fault tolerant circuits has also been explored in [26], and the concept of fault distance (which is related to spacetime code distance) is explored in [27].
Initialization of a Floquet Code
In Section 6, we provide a detailed study of the initialization behavior of Floquet codes. Each cycle of a Floquet code involves measuring the same set of operators in a fixed order. Understanding this initialization process turns out to be fruitful in understanding how many rounds it takes to do a full round of syndrome measurements.
Our first key result (Theorem 47) shows that the set of instantaneous stabilizer generators grows monotonically over successive cycles: specifically, after measuring a given operator $m_{i}$, the set of stabilizers present at that point in cycle $j$ is always a subset of those in cycle $j+1$. This result is established by applying the machinery of the distance algorithm introduced earlier in the paper. The use of this algorithm provides clarity in proving the results that would otherwise require much more elaborate tracking of stabilizers.
Building on this, Theorem 48 formalizes how the number of new stabilizer generators can evolve from cycle to cycle. We show that if $k$ new generators are added in a given cycle, then at most $k$ additional generators may appear in the subsequent cycle. Moreover, for any specific measurement $m_{i}$, a growth in the number of stabilizer generators in the current cycle implies that the same growth must have been possible in the previous cycle.
To explore the limits of how many cycles it take to fully initialize, we construct an explicit example showing that this bound is tight: that is, there exist Floquet codes that require exactly n1n{-}1 cycles to initialize a code with $n$ stabilizer generators. This is accomplished via a recursive construction of measurement sequences that add exactly one new stabilizer per cycle. Figures 6 and 7 illustrate the mechanics of this recursive design, showing how existing stabilizers are transformed and extended to generate the fully initialized set after a very long number of iterations of a measurement sequence.
Crucially, this initialization behavior is tightly linked to the understanding masking and syndrome extraction. The final theorem in this section connects the number of cycles needed for full initialization with the time required to unmask all the stabilizer generators in a given instantaneous stabilizer group (ISG). We prove that after at most $k$ cycles—where $k$ is the number of cycles required to initialize the full Floquet code—all stabilizers in the ISG will be either fully unmasked or permanently masked. This connection allows us to bound the total number of required measurements to determine all unmasked stabilizers: if each cycle involves $m$ measurements and the code requires $k$ cycles to initialize, then the number of measurements needed to fully determine the unmasked part of any ISG is at most $m\cdot k$. In particular, for code families with constant initialization depth and measurements ($k=O(1)$, $m=O(1)$), the unmasking process remains efficient and scalable.
Generalized Bravyi-Koenig theorem to dynamical code
@@
(Section 7) For a quantum code with a fixed set of stabilizer generators, the problem of finding the distance and fault-tolerant gates has been previously studied from the angle of embedding them in Euclidean space, as locality is very important in practical code design. Bravyi and Terhal [28] proved an upper bound on local code distance: $d=O(L^{D-1})$. Further, Bravyi and König [29] showed that a $D$-dimensional local code can only transversally implement encoded gates from the set $\mathcal{C}^{(D)}$, with $\mathcal{C}^{(1)}$ the Pauli group and $\mathcal{C}^{(j)}$ as the $j^{\mathrm{th}}$ level of the Clifford hierarchy. Thus, we are interested in answering this question: To what extent does these locality restrictions apply to geometrically local dynamical codes?
@@
We answer the above question and also generalize Pastawski and Yoshida’s results [30] to give a no-go theorem for 2D geometrically local dynamical code to support transversal non-Clifford gates even with some amount of long-range connectivity and non-local stabilizers. This result also extends to higher dimensions. Specifically, we prove that if a logical qubit $q_{L}$ can be supported on a region sufficiently far away from the qubits with long range connectivity, then a transversal gate that implements a logical single qubit gate on $q_{L}$ must be from the $D^{\mathrm{th}}$ level Clifford hierarchy. Furthermore, if there is limited amount of long range connectivity, then the code cannot support a logical gate from a higher Clifford hierarchy.
@@ 3.1 Stabilizer code
A stabilizer code $\mathcal{Q}=\mathcal{Q}(S)$ with parameters $[[n,k,d]]$ is a code defined on the Hilbert space $\mathcal{H}$ of $n$ qubits. It has nkn-k Pauli stabilizer generators $S$ and $k$ logical qubits. The codeword space $\mathcal{C}$ is a subspace of $\mathcal{H}$, and contains the +1+1 eigenvectors for stabilizers in $S$. The non-trivial logical representatives are given by the set $N(S)\backslash\langle S\rangle$, where $N(S)$ is the normalizer of $\langle S\rangle$ defined as $N(S)=\{p\in\mathcal{P}_{n}:\langle S\rangle p=p\langle S\rangle\},\mathcal{P}_{n}$ is the $n$ qubit Pauli group, with $\mathcal{P}=\{\pm 1,\pm i,I,X,Y,Z\}$, $\mathcal{P}_{n}=\mathcal{P}^{\otimes n}$ and $\langle S\rangle$ is the group generated by $S$. $d$ is the distance, which for a stabilizer code is the smallest weight Pauli operator $L$ for which $L$ commutes with all elements of $S$ but is not itself in $\langle S\rangle$. The distance is the lowest weight of elements in $N(S)\backslash\langle S\rangle$.
@@ 3.4 Update rules
[…]
@@
[…]
@@
    If $\pm m\notin\langle\mathcal{S}\rangle$ and $m,s]=0\,\forall s\in\mathcal{S}$, then we update the set of stabilizer generators: $\mathcal{S}\to\mathcal{S}\cup\{O(m)\cdot m\}$. This assumes that $m$ is not a logical operator.
@@
[…]
@@
[…]
@@
    If $m$ anti-commutes with $L$ and commutes with $\langle S\rangle$, then $L$ is updated to $O(m)\cdot m$. The new state is a +1+1 eigenstate of $O(m)\cdot m$ instead of $L$.
@@ 3.5 Dynamical code and Subsystem code
[…]